OpenStudy (anonymous):

Can someone explain the process of solving this problem? -What is the common difference of a 43-term arithmetic sequence where the first term is -13 and the sum is 9,374?

OpenStudy (amistre64):

id start with the formula for the sum of an arith progression

OpenStudy (amistre64):

each term can be written in terms of the first term ... lots of terms in this sentence

OpenStudy (amistre64):

can you recall any formulas for these?

OpenStudy (anonymous):

no :/ i haven't learned these

OpenStudy (amistre64):

they should be found in your material; you may want to review your material so that you have at least a basic knowledge required to proceed

OpenStudy (anonymous):

it's an online course, so i only take quizzes i don't have a teacher or anything, i basically have to look up everything online.

OpenStudy (amistre64):

those formulas can be found online as well. we will need them in order to move ahead. we could develop them if need be, but thats only an option if you truely want to learn how they work ...

OpenStudy (anonymous):

so a is the first term and d is the common difference in the formulaxn = a^sub x + d(n-1)

OpenStudy (amistre64):

thats a good start, alittle rough, but good enough to work on

OpenStudy (amistre64):

the sum of an arith progression can be calculated by adding the first and last term, and multiply the result by half the number of terms that there are.

OpenStudy (amistre64):

a1 + a1 + d(n-1) = 2a1 + d(n-1) we know there are 43 terms and that it has to equal that given amount

OpenStudy (anonymous):

okay, how to i find the last term then?

OpenStudy (amistre64):

the last term is not that important to know since it can be defined by the first term

OpenStudy (anonymous):

all i have is-13x+d(n-1) so I'm confused how to apply the formula to get what you have.

OpenStudy (amistre64):

the first term is:\[a_1\]the last terms is: \[a_n=a_1+d(n-1)\] adding these together we get\[a_1+a_1+d(n-1)=\color{red}{2a_1+d(n-1)}\]

OpenStudy (anonymous):

ahhhhh okay!

OpenStudy (amistre64):

therefore the sum is defined to be:\[\frac n2(first+last)\] \[9374=\frac{43~(\color{red}{2(-13)+d(43-1)})}{2}\] we now everything in that except for d,

OpenStudy (amistre64):

im sure you know how to do the algebra for d right ...

OpenStudy (anonymous):

yep! ill let you know what i get!

OpenStudy (anonymous):

i got -8.71 , but my answer choices are 1) 11 2) 11.5 3) 12 4) 12.5

OpenStudy (anonymous):

i didn't dived by 2 ahh hold on

OpenStudy (amistre64):

lol, lets unravel the d\[9374=\frac{43~(\color{red}{2(-13)+d(43-1)})}{2}\] \[2(9374)=43~(\color{red}{-2(13)+d(42)})\] \[\frac{2(9374)}{43}=\color{red}{-2(13)+d(42)}\] \[\frac{2(9374)}{43}+2(13)=\color{red}{d(42)}\] \[\frac{\frac{2(9374)}{43}+2(13)}{42}=\color{red}{d}\]

OpenStudy (anonymous):

but then i got -17.42 :/

OpenStudy (amistre64):

that will result in one of the options ....

OpenStudy (amistre64):

or if the mathing is tricking you up; we know the formula; so plug in the options to see what fits

OpenStudy (amistre64):

does... \[9374=\frac{43~(\color{red}{2(-13)+\color{green}{11}(43-1)})}{2}\] or \[9374=\frac{43~(\color{red}{2(-13)+\color{green}{11.5}(43-1)})}{2}\] or etc ...

OpenStudy (anonymous):

9.76. WRONG AGAIN. sorry i am very very bad with math.

OpenStudy (amistre64):

:) if your using a calculator, then its prolly trying to interpret your input keys ... 9374=43(2(-13)+d(43-1))/2 http://www.wolframalpha.com/input/?i=9374%3D43%282%28-13%29%2Bd%2843-1%29%29%2F2

OpenStudy (anonymous):

AH WOW....thank you that makes a lot of sense.

OpenStudy (amistre64):

youre welcome, and good luck :)